# Born-Infeld Electrogravity and Dyonic Black Holes

### v1

@@ V Dyonic Black Holes
Equations (53) and (60) make up an Einstein-Born-Infeld system that is equivalent to the system formed by Eqs. (37) and (40). Its spherically symmetric solutions have been widely studied, and we refer the reader to [10] (and references therein) for a comprehensive review on this topic. This geometry has been studied in standard Einstein-Born-Infeld contexts [26, 27] and also in the anomalous determinantal context [15].
We are interested in the capabilities of this type of theory for smoothing singularities; as an example we solve the field equations for the solution corresponding to a dyonic black hole, where both the electric and magnetic monopoles are present.
@@ V.1 Electrodynamics
If $\beta$ is real, then $e(r)$ is real for $r>r_{o}\equiv\left[\beta^{2}\lambda^{-2}(q^{2}+p^{2})\right]^{1/4}$, due to the anomalous behavior in the picture $g$. The magnetic field (63) is regular in the domain $r\in(r_{o},\infty)$, but $e(r)$ diverges at $r=r_{o}$, Instead, if $\beta$ is imaginary the solution is defined in the domain $r\in(0,\infty)$; the electric field is regular everywhere (it is the standard Born-Infeld field of a point-like charge), but the magnetic field diverges at $r=0$.
@@ V.2 Gravitodynamics
Then, to solve the gravitodynamic equation (60) with the source (65), $f$ must be a solution to the first order equation
@@
So, it is convenient to introduce the nondimensional coordinate $\rho\equiv r/r_{o}>1$, and the nondimensional parameter
$\alpha\equiv\frac{\epsilon}{r_{o}^{2}}\ .$ (69)
Thus, the solution to Eq. (68) reads
$f(\rho)=1-\frac{C}{\rho}-\frac{\rho^{2}}{3\ \alpha}+\frac{\lambda}{3\alpha}\ \sqrt{\rho^{4}-1}+\frac{2\lambda}{3\alpha\ \rho}\,\ \Phi(\rho)\ .$ (70)
Here $C$ is an integration constant, and $\Phi(\rho)$ is
$\Phi(\rho)\equiv 2i\ \text{K}(-1)+\text{K}(-2)-i\ \text{F}(\arcsin\rho|-1)\ ,$ (71)
where $\text{F}(\varphi|m)$ is the elliptic integral of the first kind, and $\text{K}(m)$ is the complete elliptic integral of the first kind. $\Phi(\rho)$ is a real function in the domain $\rho>1$. The constants associated with K ensure that $\Phi(\infty)=0$; so $C$ is proportional to the mass $M$ of the black hole. Thus, the behavior of $f(r)$ at infinity is
$f(r)=1-\frac{2MG}{r}-\frac{\Lambda}{3}\ r^{2}+\frac{\beta^{2}(q^{2}+p^{2})}{2\lambda\epsilon}\ r^{-2}+\frac{\beta^{4}(q^{2}+p^{2})^{2}}{40\lambda^{3}\epsilon}\ r^{-6}+\mathcal{O}(r^{-10})\ .$ (72)
@@
$\left(\frac{r}{r_{o}}\right)^{2}+\lambda\ \sqrt{\left(\frac{r}{r_{o}}\right)^{4}+1}+\frac{\epsilon}{r_{o}^{2}}\left[-1+f(r)+rf^{\prime}(r)\right]=0\ ,$ (73)
whose solution is
$f(\rho)=1-\frac{C}{\rho}+\frac{\rho^{2}}{3\ \alpha}-\frac{\lambda}{3\alpha}\ \sqrt{\rho^{4}+1}+\frac{2\lambda}{3\alpha\ \rho}\,\ \Psi(\rho)\ ,$ (74)
where
$\Psi(\rho)\equiv\sqrt{i}\left[2\ \text{K}(-1)-i\ \text{K}(2)+\ \text{F}(i\ \sinh^{-1}(\sqrt{i}\ \rho)|-1)\right]$ (75)
is a real function over the entire domain $\rho>0$ that satisfies $\Psi(\infty)=0$. The behavior of $f(r)$ at infinity is
$f(r)=1-\frac{2MG}{r}-\frac{\Lambda}{3}\ r^{2}+\frac{\beta^{2}(q^{2}+p^{2})}{2\lambda\epsilon}\ r^{-2}-\frac{\beta^{4}(q^{2}+p^{2})^{2}}{40\lambda^{3}\epsilon}\ r^{-6}+\mathcal{O}(r^{-10})\ .$ (76)
Comparing the metrics (72) and (76), we see that the choice of an imaginary parameter $\beta$ in the Lagrangian changes the metric just to the order $r^{-6}$. Compared with the Reissner-Nordström metric, which is a solution to the Einstein-Maxwell equations,
$f_{RN}(r)=1-\frac{2MG}{r}-\frac{\Lambda}{3}\ r^{2}+GQ^{2}\ r^{-2}\ ,$ (77)
we find that the role of $G$ is played by $\beta^{2}/(2\lambda\epsilon)$, as in Eq. (60). Thus, the parameter $\alpha$ defined in Eq. (69) can be rewritten as
$\alpha=\frac{\beta}{2G\sqrt{q^{2}+p^{2}}}\ .$ (78)
V.3 Metric horizons
@@
To locate the horizons of the metric, one must find the roots of the function $f(\rho)=g_{00}$. A graphic analysis of $f(\rho)$ in Eqs. (70) and (74) easily shows that there may be two, one or no horizons depending on the relation between the integration constant $C$ and the parameter $\alpha$. This is analogous to the case of Reissner-Nordström geometry, where the limit between one and two horizons is marked by the equality of the roots of $f$ when $GM^{2}=Q^{2}$ (extremal black hole). In a more general case, when roots cannot be easily computed, the extremal case can be obtained by looking for the conditions for the vanishing of $f$ and $f^{\prime}$ at the same point $\rho_{h}$. This double condition results in two equations to find $\rho_{h}$ and the relation $C=C(\alpha)$. For instance, in Reissner-Nordström one obtains
$1-\frac{2MG}{r_{h}}+\frac{GQ^{2}}{r_{h}^{2}}=0\ ,\ \ \ \ \frac{2MG}{r_{h}^{2}}-2\frac{GQ^{2}}{r_{h}^{3}}=0\ \ \ \ \Rightarrow\ \ \ r_{h}=GM\ ,\ \ \ \sqrt{G}M=Q\ ;$ (79)
at the level of nondimensional variables, this corresponds to $\rho_{h}=C/2$, $C=\sqrt{2/\alpha}$ (i.e., $\rho_{h}=1/\sqrt{2\alpha}$).
Applying this criterion to the function (70), it is found that
@@
Note that $\rho_{h}\geq 1$, so the horizon of the extremal black hole always lies inside the domain of the coordinate $\rho$ ($\rho_{h}=1$ for $\alpha=1$). The function $\rho_{h}(\alpha)$ has two branches that span the range $1<\rho_{h}<\infty$. We are interested in the branch where $0<\alpha<1$, since the Reissner-Nordström horizon is recovered in the limit $\alpha\rightarrow 0$. The relation between $C$ and $\alpha$ allowing the black hole to be extremal with horizon at $\rho_{h}$ can be obtained from $f(\rho_{h})=0$; it yields
@@
$C_{\text{extr}}=\frac{2\,\sqrt{1+\alpha^{2}}}{3\,\sqrt{2\alpha}}+\frac{2}{3\,\alpha\ }\,\ \Phi\left(\frac{\sqrt{1+\alpha^{2}}}{\sqrt{2\alpha}}\right)\ .$ (81)
Instead, if the determinantal Lagrangian is built with an imaginary value of $\beta$, then the geometry will be characterized by $f(\rho)$ in Eq. (74). By applying the same procedure, one obtains the horizon of the extremal black hole at
$\rho_{h}=\frac{\sqrt{1-\alpha^{2}}}{\sqrt{2\alpha}}\ ,$ (82)
while the relation between $C$ and $\alpha$ becomes
@@
$C_{\text{extr}}=\frac{2\,\sqrt{1-\alpha^{2}}}{3\,\sqrt{2\alpha}}+\frac{2}{3\,\alpha\ }\,\ \Psi\left(\frac{\sqrt{1-\alpha^{2}}}{\sqrt{2\alpha}}\right)\ .$ (83)
In both cases the Reissner-Nordström limit is retrieved:
@@
$C_{\text{extr}}=\sqrt{\frac{2}{\alpha}}+\mathcal{O}(\alpha^{3/2})\ .$ (84)
Figure 1 shows how the existence and number of horizons depend on the relation between the parameters $\alpha$ and $C$.
@@
Figure 1: Plot of the function $f(\rho)$ when $C=C_{\text{extr}}$ (Eq. (70)) and when small deviations of that value are considered. It was chosen $\lambda=1$ and $\alpha=0.1$.
For $C$ (the parameter proportional to the black-hole mass) greater than $C_{\text{extr}}$ the curve exhibits two roots, which indicates the presence of two horizons. For $C<C_{\text{extr}}$, there are no horizons. A similar graphic is obtained for the solution (74) associated with an imaginary value of $\beta$ in the determinantal Lagrangian (although $\rho\in(0,\infty)$, and $f$ diverges at $\rho=0$).
V.4 The smoothing of singularities

### v2

@@ V Dyonic Black Holes
Equations (53) and (60) make up an Einstein–Born–Infeld system that is equivalent to the system formed by Eqs. (37) and (40). Its spherically symmetric solutions have been widely studied, and we refer the reader to [10] (and references therein) for a comprehensive review on this topic. This geometry has been studied in standard Einstein–Born–Infeld contexts [31, 32] and also in the anomalous determinantal context [15].
We are interested in the capabilities of this type of theory for smoothing singularities; as an example we solve the field equations for the solution corresponding to a dyonic BH, where both the electric and magnetic monopoles are present.
@@ V.1 Electrodynamics
If $\beta$ is real in the Lagrangian (i.e., picture $g$ is anomalous), then $e(r)$ is real only for $r>r_{o}\equiv\left[\beta^{2}\lambda^{-2}(q^{2}+p^{2})\right]^{1/4}$. The magnetic field (63) is regular in the domain $r\in(r_{o},\infty)$, but $e(r)$ diverges at $r=r_{o}$, Instead, if $\beta$ is imaginary in the Lagrangian (i.e., picture $g$ is standard), then the solution is defined in the domain $r\in(0,\infty)$; the electric field is regular everywhere (it is the standard BI field of a point-like charge), but the magnetic field diverges at $r=0$.
@@ V.2 Gravitodynamics
To solve the gravitodynamic equation (60) with the source (65), $f$ must be a solution to the first order equation
@@
It will be convenient to switch to nondimensional coordinate and parameter:
$\rho\equiv\frac{r}{r_{o}}=\frac{\sqrt{\lambda\beta^{-1}}\ r}{(q^{2}+p^{2})^{1/4}}\ ,\ \ \ \ \ \alpha\equiv\frac{\epsilon}{r_{o}^{2}}=\frac{\beta}{2G\sqrt{q^{2}+p^{2}}}\ ,$ (69)
where the relationship $G=\beta^{2}/(2\lambda\epsilon)$ emerges from the Eq. (60). Thus, the solution to Eq. (68) is defined in the domain $\rho\geq 1$ and turns out to be
$f(\rho)=1-\frac{C}{\rho}-\frac{\rho^{2}}{3\ \alpha}+\frac{\lambda}{3\alpha}\ \sqrt{\rho^{4}-1}+\frac{2\lambda}{3\alpha\ \rho^{2}}\,\ {{}_{2}F_{1}}\left(\frac{1}{4},\frac{1}{2};\frac{5}{4};\rho^{-4}\right)\ .$ (70)
Here, ${{}_{2}F_{1}}(\frac{1}{4},\frac{1}{2};\frac{5}{4};\frac{1}{x})$ is the ordinary hypergeometric function, which is real and finite for $x\in(-\infty,0]\cup[1,\infty)$. $C$ is an integration constant proportional to the BH mass $M$, $C=2MGr_{o}^{-1}$. In Figure 1(a) we display the function $f(\rho)$ for different values of CC. [^9] For further checks, the hypergeometric function satisfies ${{}_{2}F_{1}}(\frac{1}{4},\frac{1}{2};\frac{5}{4};\rho^{-4})=\rho\,[2i\ \text{K}(-1)+\text{K}(2)-i\ \text{F}(\arcsin\rho|-1)]$, where $\text{F}(\varphi|m)$ is the elliptic integral of the first kind, and $\text{K}(m)$ is the complete elliptic integral of the first kind. In addition, ${{}_{2}F_{1}}(\frac{1}{4},\frac{1}{2};\frac{5}{4};-\rho^{-4})=\rho\,\sqrt{i}\left[2\ \text{K}(-1)-i\ \text{K}(2)+\ \text{F}(i\ \sinh^{-1}(\sqrt{i}\ \rho)|-1)\right]$.
The behavior of $f(r)$ at infinity is
$f(r)=1-\frac{2MG}{r}-\frac{\Lambda}{3}\ r^{2}+\frac{G\,(q^{2}+p^{2})}{r^{2}}+\frac{\beta^{2}G(q^{2}+p^{2})^{2}}{20\lambda^{2}\,r^{6}}+\mathcal{O}(r^{-10})\ .$ (71)
@@
$\left(\frac{r}{r_{o}}\right)^{2}+\lambda\ \sqrt{\left(\frac{r}{r_{o}}\right)^{4}+1}+\frac{\epsilon}{r_{o}^{2}}\left[-1+f(r)+rf^{\prime}(r)\right]=0\ .$ (72)
This equation admits the solution
$f(\rho)=1-\frac{C}{\rho}+\frac{\rho^{2}}{3\ \alpha}-\frac{\lambda}{3\alpha}\ \sqrt{\rho^{4}+1}+\frac{2\lambda}{3\alpha\ \rho^{2}}\,\ {{}_{2}F_{1}}\left(\frac{1}{4},\frac{1}{2};\frac{5}{4};-\rho^{-4}\right)\ ,$ (73)
which is a real function throughout the domain $\rho>0$. A plot analogous to Figure  1(a) is obtained in this case, although here the domain extends to $\rho\in(0,\infty)$ and the function diverges as $\rho\to 0$. The behavior of $f(r)$ at infinity is
$f(r)=1-\frac{2MG}{r}-\frac{\Lambda}{3}\ r^{2}+\frac{G\,(q^{2}+p^{2})}{r^{2}}-\frac{\beta^{2}G(q^{2}+p^{2})^{2}}{20\lambda^{2}\,r^{6}}+\mathcal{O}(r^{-10})\ .$ (74)
Comparing the metrics (71) and (74), we see that the choice of an imaginary parameter $\beta$ in the Lagrangian changes the metric to just the order $r^{-6}$, which is also the order in which these metrics depart from the Reissner-Nordström metric
$f_{RN}(r)=1-\frac{2MG}{r}-\frac{\Lambda}{3}\ r^{2}+\frac{G\,Q^{2}}{r^{2}}\ .$ (75)
V.3 Number of horizons. Extremal BH
@@
To locate the horizons of the metric, one looks for the roots of the function $f(\rho)=g_{00}$. A graphic analysis of $f(\rho)$ in Eqs. (70) and (73) easily shows that there may be two, one, or no horizons depending on the relation between the integration constant $C$ and the parameter $\alpha$. This is analogous to the case of Reissner-Nordström geometry, where the limit between one and two horizons is marked by the equality of the roots of $f$ when $GM^{2}=Q^{2}$ (extremal BH). In a more general case, when roots cannot be easily computed, the extremal case can be obtained by looking for the conditions for the vanishing of $f$ and $f^{\prime}$ at the same point $\rho_{h}$. This double condition results in two equations to find $\rho_{h}$ and the relation $C=C(\alpha)$. For instance, in Reissner-Nordström one obtains
$1-\frac{2MG}{r_{h}}+\frac{GQ^{2}}{r_{h}^{2}}=0\ ,\ \ \ \ \frac{2MG}{r_{h}^{2}}-2\frac{GQ^{2}}{r_{h}^{3}}=0\ \ \ \ \Rightarrow\ \ \ r_{h}=GM\ ,\ \ \ \ \sqrt{G}M=|Q|\ .$ (76)
At the level of nondimensional variables, this result corresponds to $\rho_{h}=C/2$, $C=\sqrt{2/\alpha}$ (i.e., $\rho_{h}=1/\sqrt{2\alpha}$).
Applying this criterion to the function (70) i.e., the solution with real $\beta$ in the Lagrangian), it is found that
@@
Note that $\rho_{h}\geq 1$; i.e., the horizon of the extremal BH always lies within the domain of the coordinate $\rho$. The function $\rho_{h}(\alpha)$ has two branches spanning the range $1<\rho_{h}<\infty$. They are related by the equality $\rho_{h}(\alpha)=\rho_{h}(\alpha^{-1})$.Thus, both branches meet at $\alpha=1$ where $\rho_{h}=1$. Therefore, for each value of $\rho_{h}$ there are two possible charges $\sqrt{q^{2}+p^{2}}$ realizing that value. The relation between $C$ and $\alpha$ allowing the BH to be extremal with horizon at $\rho_{h}$ can be obtained from $f(\rho_{h})=0$; it yields
@@
$C_{\text{extr}}$ $=$ $\frac{\sqrt{2}\,\sqrt{1+\alpha^{2}}}{3\,\sqrt{\alpha}}\left[1+\frac{2}{1+\alpha^{2}}\,\ {{}_{2}F_{1}}\left(\frac{1}{4},\frac{1}{2};\frac{5}{4};\left(\frac{2\alpha}{1+\alpha^{2}}\right)^{2}\right)\right]\ ,\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \mathrm{if}\ \ \ 0<\alpha<1\,$, (78)
@@
$C_{\text{extr}}$ $=$ $\frac{\sqrt{2}\,\sqrt{1+\alpha^{2}}}{3\,\sqrt{\alpha}}\left[1-\frac{1-\alpha^{4}}{2\alpha^{2}\,(1+\alpha^{2})}+\frac{2}{1+\alpha^{2}}\,\ {{}_{2}F_{1}}\left(\frac{1}{4},\frac{1}{2};\frac{5}{4};\left(\frac{2\alpha}{1+\alpha^{2}}\right)^{2}\right)\right]\ ,\ \ \ \mathrm{if}\ \ \ \alpha>1\,.$ (79)
The extremal BH mass is
$M_{\text{extr}}=\frac{r_{o}}{2G}\ C_{\text{extr}}=\frac{\sqrt{\epsilon}}{2G\sqrt{\alpha}}\ C_{\text{extr}}=\frac{\beta}{(2G)^{3/2}\sqrt{\alpha}}\ C_{\text{extr}}\ .$ (80)
Given that $\alpha$ is a measure of the inverse of the charge in terms of the fundamental unit of charge $\beta/G$ (see Eq. (69)), then the relationship between mass and charge for the extremal BH is contained in the previous equations: Figure 1(b) shows the curve $\sqrt{G}M_{\text{extr}}$ vs. $\alpha$. At the point $\alpha=1$ where the two above mentioned branches meet, it is
@@
$\sqrt{G}M_{\text{extr}}(\alpha=1)=\frac{\beta}{2G}\,\frac{2}{3}\left[1+\sqrt{\pi}\ \frac{\Gamma(\frac{5}{4})}{\Gamma(\frac{3}{4})}\right]=\frac{2}{3}\sqrt{q^{2}+p^{2}}\,\left[1+\sqrt{\pi}\ \frac{\Gamma(\frac{5}{4})}{\Gamma(\frac{3}{4})}\right]=1.54\,\sqrt{q^{2}+p^{2}}\ .$ (81)
(we applied Eq. (69) with $\alpha=1$). Besides,
$\sqrt{G}M_{\text{extr}}(\alpha)$ $\xrightarrow[\alpha\to 0$]{} $\frac{\beta}{4\,G}\,\left(\frac{2}{\alpha}+\frac{\alpha}{5}+{\cal O}(\alpha^{3})\right)=\sqrt{q^{2}+p^{2}}+\frac{\beta^{2}}{40\,G^{2}\sqrt{q^{2}+p^{2}}}+{\cal O}(\alpha^{3})\ ,$ (82)
$\sqrt{G}M_{\text{extr}}(\alpha)$ $\xrightarrow[\alpha\to\infty$]{} $\frac{\beta}{4\,G}+\frac{3\beta}{8\alpha^{2}}+{\cal O}(\alpha^{-3})=\frac{\beta}{4\,G}+\frac{3\,G^{2}}{2\beta}\,(q^{2}+p^{2})+{\cal O}(\alpha^{-3})\ .$ (83)
Very interestingly, in BI electrogravity the limit for small charge ($\alpha\rightarrow\infty$, $\beta$ remains finite) leads to an extremal BH of fundamental mass $M_{\text{extr}}=\beta/(4\,G^{3/2})$. The horizon of the fundamental extremal BH is located at
$r_{h}=r_{o}\,\rho_{h}\xrightarrow[\alpha\to\infty]{}r_{o}\,\sqrt{\frac{\alpha}{2}}=\sqrt{\frac{\epsilon}{2}}=\frac{\beta}{2\,\sqrt{G}}\ .$ (84)
Instead, if the determinantal Lagrangian is built with an imaginary value of $\beta$, then the geometry will be characterized by $f(\rho)$ in Eq. (73). By applying the same procedure, one obtains the horizon of the extremal BH at
$\rho_{h}=\frac{\sqrt{1-\alpha^{2}}}{\sqrt{2\alpha}}$ (85)
($0<\alpha<1$), while the relation between $C$ and $\alpha$ becomes
$C_{\text{extr}}=\frac{\sqrt{2}\,\sqrt{1-\alpha^{2}}}{3\,\sqrt{\alpha}}\left[1+\frac{2}{1-\alpha^{2}}\,\ {{}_{2}F_{1}}\left(\frac{1}{4},\frac{1}{2};\frac{5}{4};-\left(\frac{2\alpha}{1-\alpha^{2}}\right)^{2}\right)\right]\ .$ (86)
In both cases the extremal Reissner-Nordström BH is retrieved in the Maxwellian limit $\beta\rightarrow 0$ (i.e., $\alpha\rightarrow 0$). In fact, it is ${{}_{2}F_{1}}\left(\frac{1}{4},\frac{1}{2};\frac{5}{4};0\right)=1$; then
$C_{\text{extr}}\xrightarrow[\beta\to 0]{}\sqrt{\frac{2}{\alpha}}\ ,\ \ \ \ \ \ \rho_{h}\xrightarrow[\beta\to 0]{}\frac{1}{\sqrt{2\alpha}}\ .$ (87)
@@
Figure 1: Left: Plot of the function $f(\rho)$ when $C=C_{\text{extr}}$ (Eq. (70)) and for small deviations from that value. We chose $\lambda=1$ and $\alpha=0.1$. Right: Extremal mass $\sqrt{G}\,M_{\text{extr}}$ as a function of $\alpha$, showing the matching of the two branches at $\alpha=1$, the RN-extremal behavior for small $\alpha$, and the asymptotic constant behavior for large $\alpha$.
V.4 Thermodynamics
The BH temperature is $T=\kappa/(2\pi)$ where $\kappa$ is the horizon surface gravity. For metrics of the form (62), the surface gravity can be computed as $\kappa=(1/2)\,\partial f/\partial r$ (for example, see [33]). The derivative must be specialized in the exterior horizon $r_{+}$, where $f(r_{+})=0$. The extremal BH has $T=0$ because $\partial f/\partial r$ is zero in $r=r_{h}$.
For the geometry (70), with $\lambda=1$, it results
$T=\frac{1}{4\pi}\frac{\partial f}{\partial r}|_{r=r_{+}}=\frac{1}{4\pi r_{o}}\frac{\partial f}{\partial\rho}|_{\rho=\rho_{+}}=\frac{1}{8\pi r_{o}}\frac{\alpha-\rho_{+}^{2}+\sqrt{\rho_{+}^{4}-1}}{\alpha\,\rho_{+}}\ .$ (88)
Although $\rho_{+}$ cannot be analitically solved, the positivity of $T$ is guaranteed, since the derivative of $f(\rho)$ at the exterior horizon (see the solutions with two horizons in Figure 1(a)). On the other hand, $f(r_{+})=0$ implies the following relation between the mass $M=Cr_{o}/(2G)$ and $\rho_{+}$:
$M=\frac{r_{o}}{6G\alpha}\left[-\rho_{+}^{3}+3\alpha\rho_{+}+\rho_{+}\sqrt{\rho_{+}^{4}-1}+2\ {{}_{2}F_{1}}\left(\frac{1}{4},\frac{1}{2};\frac{5}{4};\rho^{-4}\right)\right]\ .$ (89)
If $q^{2}+p^{2}$ goes to zero, then $\alpha$ goes to infinity (see Eq. (69)); then the Schwarzschild relation $T=1/(4\pi GM)$ is obtained from Eqs. (88) and (89).
Figure 2: Parametric plot of the relation $T$ vs $M$, obtained from Eqs. (88) and (89).
The general relation $T$ vs. $M$ can be visualized through a parametric plot, starting from Eqs. (88) and (89). Figure 2 shows that the specific heat $\partial M/\partial T$ is positive for small masses and negative for large masses; it diverges at the value of $M$ where the temperature reaches its maximum. This behaviour is typical of charged BHs.
The geometry (73) has already been considered in the literature, since it is sourced by the standard BI energy-momentum tensor of a point-like charge. Its thermodynamics has also been studied, although in somewhat different contexts [34, 35, 36, 37]. In this case, the results are as follows,
$T=\frac{1}{8\pi r_{o}}\frac{\alpha+\rho_{+}^{2}-\sqrt{\rho_{+}^{4}+1}}{\alpha\,\rho_{+}}\ ,$ (90)
$M=\frac{r_{o}}{6G\alpha}\left[\rho_{+}^{3}+3\alpha\rho_{+}-\rho_{+}\sqrt{\rho_{+}^{4}+1}+2\ {{}_{2}F_{1}}\left(\frac{1}{4},\frac{1}{2};\frac{5}{4};-\rho^{-4}\right)\right]\ .$ (91)
The graph $T$ vs. $M$ exhibits similar characteristics than the previous one.
V.5 The smoothing of singularities